Errata for CERTAIN MINIMAL OR HOMOLOGICALLY VOLUME MINIMIZING SUBMANIFOLDS IN COMPACT SYMMETRIC SPACES

(Tsukuba J. Math. Vol. 9 No. 1 (1985), 117–131)

Hiroyuki Tasaki

October 12, 1999

p.120 Line 2 of Lemma 1: “𝔪 = 𝔞 + ∑_{γ∈Σ_{+}} 𝔪_{γ}.” should read “𝔪 = 𝔞 + ∑_{γ∈Σ_{+}} 𝔪_{γ}”
p.123 Line 16: “impies” should read “implies”
p.130 Line 3 from bottom: “and isometry g” should read “an isometry g”
p.131 Line 3 from bottom: “affermatively” should read “affirmatively”